OpenStudy (lovely_bae):

What is -3x-2y=20 in slope intercept form

OpenStudy (phi):

I would first add +3x to both sides, like this -3x+3x -2y = 20+3x now simplify the left side. can you do that ?

OpenStudy (lovely_bae):

ok -3x+3x=6x = 23x so in other words 6x*-2y=23x.....

OpenStudy (phi):

-3x+3x means you have 3 x's and take away 3 x's how many x's do you have ?

OpenStudy (lovely_bae):

0

OpenStudy (phi):

yes, so -3x+3x -2y = 20+3x becomes -2y = 20+3x on the right side, we can change the order (when adding, order does not matter) so -2y = 3x + 20 now divide both sides by -2 (both sides all terms) on the left side -2/-2 becomes 1 so you get 1y or just y on the left side what do you get on the right side ?

OpenStudy (lovely_bae):

1.5.....

OpenStudy (lovely_bae):

?

OpenStudy (lovely_bae):

.............

OpenStudy (tgstudios):

it would be -1.5

OpenStudy (tgstudios):

so... y = -1.5x + -10 or - 10

OpenStudy (tgstudios):

so your answer would be y = -1.5 - 10 :)

OpenStudy (phi):

you get \[ y = \frac{3x}{-2}+ \frac{20}{-2} \\ y = -\frac{3}{2} x -10\]

OpenStudy (tgstudios):

orrrrrr, what I said XD

OpenStudy (tgstudios):

@lovely_bae ? did you see this?

OpenStudy (lovely_bae):

ahhh yezzsirr